a) What is the Young’s modulus of metal is the speed of the sound in a rod is 4550 m/s and it’s density is p=5230 kg/m3?

b) What is the fundamental frequency and second harmonic of a sound wave that propagates inside an open tube, if the air is at a temperature of 35 Celsius degrees? The length of the tube is 40 cm.


1
Expert's answer
2021-09-28T16:55:07-0400

(a) "v=\sqrt{E/\rho}\to E=v^2\rho=4550^2\cdot5230=108.3\ (GPa)"


(b) "v=\sqrt{\gamma\frac{RT}{M}}=\sqrt{1.4\cdot\frac{8.31\cdot308}{0.029}}=352\ (m/s)"


"f=\frac{nv}{2L}\to f_1=\frac{352}{2\cdot0.4}=440\ (Hz)"


"f_2=\frac{2\cdot352}{2\cdot0.4}=880\ (Hz)"